OpenStudy (anonymous):

Factor completely. If the polynomial is prime, state this. y2 + 0.4y + 0.04

OpenStudy (anonymous):

The expression is of the form of a perfect square:\[y^2 + 0.4y + 0.04=y^2 + 2y/5 + 1/25=(y+1/5)^2\]